A USA Today article claims that the proportion of people who believe global warming is a serious issue is 0.73, but given the number of people you've talked to about this same issue, you believe it is different from 0.73. The hypotheses for this test are Null Hypothesis: p = 0.73, Alternative Hypothesis: p ≠ 0.73. If you randomly sample 21 people and 12 of them believe that global warming is a serious issue, what is your test statistic and p-value?

Question 7 options:

1)

Test Statistic: 1.637, P-Value: 0.102

2)

Test Statistic: -1.637, P-Value: 0.898

3)

Test Statistic: -1.637, P-Value: 0.051

4)

Test Statistic: -1.637, P-Value: 0.949

5)

Test Statistic: -1.637, P-Value: 0.102

Question 8 (1 point)

Your friend tells you that the proportion of active Major League Baseball players who have a batting average greater than .300 is less than 0.8, a claim you would like to test. The hypotheses for this test are Null Hypothesis: p ≥ 0.8, Alternative Hypothesis: p < 0.8. If you randomly sample 20 players and determine that 13 of them have a batting average higher than .300, what is the test statistic and p-value?

Question 8 options:

1)

Test Statistic: 1.677, P-Value: 0.953

2)

Test Statistic: -1.677, P-Value: 0.094

3)

Test Statistic: -1.677, P-Value: 0.953

4)

Test Statistic: 1.677, P-Value: 0.047

5)

Test Statistic: -1.677, P-Value: 0.047

Question 9 (1 point)

As of 2012, the proportion of students who use a MacBook as their primary computer is 0.33. You believe that at your university the proportion is actually less than 0.33. The hypotheses for this scenario are Null Hypothesis: p ≥ 0.33, Alternative Hypothesis: p < 0.33. You conduct a random sample and run a hypothesis test yielding a p-value of 0.7816. What is the appropriate conclusion? Conclude at the 5% level of significance.

Question 9 options:

1)

The proportion of students that use a MacBook as their primary computer is significantly less than 0.33.

2)

We did not find enough evidence to say a significant difference exists between the proportion of students that use a MacBook as their primary computer and 0.33

3)

We did not find enough evidence to say the proportion of students that use a MacBook as their primary computer is larger than 0.33.

4)

The proportion of students that use a MacBook as their primary computer is greater than or equal to 0.33.

5)

We did not find enough evidence to say the proportion of students that use a MacBook as their primary computer is less than 0.33.

Question 10 (1 point)

As of 2012, the proportion of students who use a MacBook as their primary computer is 0.42. You believe that at your university the proportion is actually different from 0.42. The hypotheses for this scenario are Null Hypothesis: p = 0.42, Alternative Hypothesis: p ≠ 0.42. You conduct a random sample and run a hypothesis test yielding a p-value of 0.7706. What is the appropriate conclusion? Conclude at the 5% level of significance.

Question 10 options:

1)

We did not find enough evidence to say a significant difference exists between the proportion of students that use a MacBook as their primary computer and 0.42

2)

The proportion of students that use a MacBook as their primary computer is significantly different from 0.42.

3)

The proportion of students that use a MacBook as their primary computer is equal to 0.42.

4)

We did not find enough evidence to say the proportion of students that use a MacBook as their primary computer is less than 0.42.

5)

We did not find enough evidence to say the proportion of students that use a MacBook as their primary computer is larger than 0.42.
Answer #1

Question 7:

Correct option:

(5)   Test Statistic: -1.637, P-Value: 0.102

Explanation:

H0: P = 0.73

HA: P 0.73

= 12/21 = 0.5714

Test Statistic is given by:

df = 21 - 1 = 20

By Technology, p value = 0.1017

Question 8:

Correct option:

(5) Test Statistic: -1.677, P-Value: 0.047

Explanation:

H0: P 0.8

HA: P < 0.8

= 13/20=0.65

Test Statistic is given by:

df = 20 - 1 = 19

By Technology, p - value = 0.0468

Question 9:

Correct option:

We did not find enough evidence to say the proportion of students that use a MacBook as their primary computer is less than 0.33.

Explanation:

H0: P 0.33

HA: P < 0.33

Since p - value = 0.7816 is greater than = 0.05, the difference is not significant. Fail to reject null hypothesis.

Question 10:

(1)   We did not find enough evidence to say a significant difference exists between the proportion of students that use a MacBook as their primary computer and 0.42

Explanation:
H0: P = 0.42

HA: P 0.42

Since p - value = 0.7706 is greater than = 0.05, the difference is not significant. Fail to reject null hypothesis.